surface patch for S. regular surface and f: S Ra smooth EXERCISE 3.44. Let S be a function. Assume that the point p e S is a critical point of f, which means that dfp(v) 0 for all v e TpS. Define the Hessian of f atp in the direction v as Hess(f)p(v) (foy)"(0), where y is a regular curve in S with y(0) = p and y'(0) = v. Prove that the Hessian is well defined in the sense that it is independent of the choice of HINT: Work in local coordinates and use Exercise 3.5 on page 123.
a diifeon EXERCISE 3.5. Let U C R2 be open andf : U Rsmooth. Denote the coordinates of R2 by fu, v}. Assume that the point q e U is a critical point of f, which means that dfg(w) = 0 for all wE R2, or equivalently, fu(a)= fu(a) =0. Let y: I U be a regular curve with (0) = q and (0)= = (a, b). Prove that (fo7'(0)=a2 fun(4) +2abfue(g) + 62for (9) a2 fun(g) +2abfuv (g) +b fr(g)
Answer #1

S a regular curve and R f:S- a smooth function. pES is a critical point of f. The Hessian is defined by

Hess(f(u) (f o(0) where pES,UETS and \gamma is a regular curve with (0) p,(0)= v .

Suppose \widetilde{\gamma} is another such curve, we have to show that o(0)fo (0) .

Let U \subset \mathbb{R}^2, q \in U and O:U S be a regular surface path such that \sigma (q)=p. Now f \circ \sigma : U \rightarrow \mathbb{R} is a smooth function with q as a critical point (that is because p is a critical point and \sigma is a diffeomorphism).

The two curves \gamma and \widetilde{\gamma} decend to a curve in U, call them \alpha and \widetilde{\alpha} , such that \sigma \circ \alpha = \gamma, \sigma \circ \widetilde{\alpha} = \widetilde{\gamma} .

According to exercise 3.5, (f \circ \sigma \circ \alpha)^{''}(0) = a^2 (f \circ \sigma)_{uu} (q) + 2ab (f \circ \sigma)_{uv} (q) + b^2 (f \circ \sigma)_{vv} (q) , where a and b depend only on v.

Since the right hand side of the equation does not depend on \alpha, we have that it is the same for \widetilde{\alpha} . Hence (f \circ \sigma \circ \alpha)^{''}(0) = (f \circ \sigma \circ \widetilde{\alpha})^{''}(0)

\implies (f \circ \gamma)^{''}(0) = (f \circ \widetilde{\gamma})^{''}(0)

Therefore the Hessian is well defined.